OpenStudy (luigi0210):

Integrate:

OpenStudy (luigi0210):

\[\LARGE \int~\sqrt{x^2+1}~dx\]

OpenStudy (raden):

try this x = tan(θ)

OpenStudy (luigi0210):

Change it to \(\Large \int~\sqrt{tan^2u+1}du\)? Then use the trig identity?

OpenStudy (luigi0210):

Whoops, \(*sec^2u\) on the outside.

OpenStudy (rational):

that gives u sec^3 and u can spend an hour or two integrating it.. if you're allowed to use tables, u may pick it directly from the table (7th integral) http://integral-table.com/downloads/single-page-integral-table.pdf

OpenStudy (anonymous):

who allows the table x-(

OpenStudy (rational):

ikr... :)

OpenStudy (anonymous):

well i would say , does this work ? 

OpenStudy (anonymous):

\[x=\tan \theta~~~dx = \sec^2 \theta d \theta \] \[\sqrt{x^2+1} = \sqrt{\tan^2 \theta +1}\] \[\int\limits \sec^3 \theta d \theta \] and now you can use a reduction formula.

OpenStudy (rational):

or an hyperbolic substitution again gives u the answer fast

random231 (random231):

ahem there is a direct formula tho: http://prntscr.com/3qfghx

OpenStudy (isaiah.feynman):